OpenStudy (anonymous):

Simplify the expression (2x − 9)(x + 6).

OpenStudy (anonymous):

given answers?

OpenStudy (anonymous):

@Brinannuh

OpenStudy (whpalmer4):

Multiply the expression out and simplify by collecting like terms.

OpenStudy (anonymous):

No multiple choice, I'm stuck. @fathomboy

OpenStudy (anonymous):

Ill expand it out but I will have you do the simplifying and calculations

OpenStudy (whpalmer4):

\[(2x-9)(x+6)=2x(x+6)-9((x+6)\]by the distributive property of multiplication. Now do that again to remove the remaining parentheses and collect like terms to simplify.

OpenStudy (anonymous):

nvm, looks like whpalmer4's got it under control

OpenStudy (whpalmer4):

Example: \[(2x-5)(x+1)=2x(x+1)-5(x+1)\]\[\qquad=2x*x+2x*1-5*x-5*1\]\[\qquad=2x^2+2x-5x-5\]\[\qquad=2x^2-3x-5\]

OpenStudy (whpalmer4):

@Brinannuh are you still stuck?

OpenStudy (anonymous):

no, thank you! @whpalmer4